### v3

@@ 2.1 Datasets
utilizes 10 large medical datasets across diverse imaging modalities (radiology, pathology, dermatology, ophthalmology) and clinical tasks for robust evaluation (Fig. 1a). Disease diagnosis datasets include CheXpert (CXP) [irvin2019chexpert], Camelyon [bandi2018detection], HAM10000 (HAM) [tschandl2018ham10000], PAPILA [papineni2002bleu], GF3300 [luo2024harvard] , and FairVLMed [luo2024fairclip]. Medical VQA datasets included VQA-RAD [lau2018dataset], PathVQA [he2020pathvqa], SLAKE [liu2021slake], and FairVLMed [luo2024fairclip], each containing both closed-form (yes/no) and open-form question–answer pairs. FairVLMed is simultaneously used for both tasks. GPT-4 is used to design one closed-form question and one open-form question from each clinical report in FairVLMed for VQA evaluation. Across all datasets, no data points were missing, and we relied exclusively on the existing data. Dataset statistics are reported in Appendix B (Tables 3 and 4) and details of the VQA dataset construction are provided in the Appendix C.
@@ 3 Results
Together, these two views allow us to address all our research questions. The complete raw data underlying these figures are included in Appendix D, with Tables 5–7 reporting diagnostic results and Tables 5–5 reporting VQA results.

### v4

@@ 2.1 Datasets
utilizes 12 large medical datasets spanning diverse imaging modalities (radiology, pathology, dermatology, ophthalmology, and multimodal clinical reasoning) and clinical tasks for robust evaluation (Fig. 2a). Disease diagnosis datasets include CheXpert (CXP) [irvin2019chexpert], Camelyon [bandi2018detection], HAM10000 (HAM) [tschandl2018ham10000], PAPILA [papila], GF3300 [luo2024harvard], and FairVLMed [luo2024fairclip]. Medical VQA datasets include VQA-RAD [lau2018dataset], PathVQA [he2020pathvqa], SLAKE [liu2021slake], OmniMedVQA [hu2024omnimedvqa], MedXpertQA [zuo2025medxpertqa], and FairVLMed [luo2024fairclip]. Among these, OmniMedVQA is a large-scale benchmark covering 12 medical imaging modalities and more than 20 anatomical regions, while MedXpertQA introduces expert-level multimodal medical reasoning questions with rich clinical context. FairVLMed is used for both diagnosis and VQA tasks. For FairVLMed, GPT-4 was used to generate one closed-form question and one open-form question from each clinical report for VQA evaluation. Across all datasets, no data points were missing, and we relied exclusively on the existing data. Dataset statistics are reported in Appendix B (Tables 3 and 4), and details of the VQA dataset construction are provided in Appendix C.
@@ 3 Results
Together, these two views allow us to address all our research questions. The complete raw data underlying these figures are included in Appendix D, with Tables 5–7 reporting diagnostic results and Tables 8–27 reporting VQA results.